### v1

@@ 5.3 Surrogate Loss Method in Both Realizable and Agnostic Cases
Again we focus empirical precision and recall losses minimization. Let $x_{1},\ldots,x_{m}\in\mathcal{X}$ denote a sequence of users. For each graph $g$, let $U_{i}^{g}$ denote the uniform distribution over the neighbors $N_{g}(x_{i})$. for any set $S\subset N_{g}(x_{i})$. For any pair of graphs $g^{\prime},g^{\prime\prime}$, define the following:
$\mathtt{precision.loss}(g^{\prime}\mid g^{\prime\prime})$ $=\mathtt{recall.loss}(g^{\prime\prime}\mid g^{\prime})=\frac{1}{m}\sum_{i=1}^{m}U_{i}^{g^{\prime}}(N_{g^{\prime}}(x_{i})\setminus N_{g^{\prime\prime}}(x_{i}))\,.$
Here $\mathtt{precision.loss}(g^{\prime}\mid g^{\prime\prime})$ is the precision loss of graph $g^{\prime}$ when the target graph is $g^{\prime\prime}$ and $\mathtt{recall.loss}(g^{\prime\prime}\mid g^{\prime})$ is the recall loss of graph $g^{\prime\prime}$ when the target graph is $g^{\prime}$. Thus, the goal is to output a graph $g$ with small $\mathtt{precision.loss}(g\mid g^{\text{target}})$ and $\mathtt{recall.loss}(g\mid g^{\text{target}})$.
@@
For a graph $g$ define a vector $v_{g}:\mathcal{H}\times\mathcal{H}\to[0,1]$ by
$v_{g}(g^{\prime},g^{\prime\prime})=\frac{1}{m}\sum_{i=1}^{m}U_{i}^{g}(N_{g^{\prime}}(x_{i})\setminus N_{g^{\prime\prime}}(x_{i})).$
Define a metric $d_{\mathcal{H}}$ between graphs by $d_{\mathcal{H}}(h,k)=\|v_{h}-v_{k}\|_{\infty}$.
Let $\widehat{g}$ be the observed (empirical) graph; i.e. the graph in which every $x_{i}$ is connected to the random number $v_{i}$ which is observed in the training set. A standard union bound argument yields:
@@
Let $g^{\text{target}}$ denote the true graph (i.e. the data is generated from $g^{\text{target}}$). Then, with probability at least $1-\delta$:
@@
Algorithm (realizable case): Let $\varepsilon$ denote the desired error. Output a graph $g^{\text{output}}\in\mathcal{H}$ such that
@@
Notice that Item 1 corresponds to the first principle for discarding suboptimal graphs described earlier in this section, while Item 2 corresponds to the second principle.
Algorithm (agnostic case): output a graph $g^{\text{output}}\in\mathcal{H}$ such that
@@
Let $g^{\text{target}}$ denote the target graph. Then, for
@@
the agnostic-case algorithm outputs a graph $g^{\text{output}}$ such that with probability at least $1-\delta$,
@@ 5.3.1 Proof of Theorem 3
Proof of Theorem 3  For simplicity, we adopt payoffs instead of losses here. The payoff of graph $g$ at $x$ is
$u(g,x)=\frac{|N_{g^{\text{target}}}(x)\cap N_{g}(x)|}{2|N_{g}(x)|}+\frac{|N_{g^{\text{target}}}(x)\cap N_{g}(x)|}{2|N_{g^{\text{target}}}(x)|}\,.$
If $N_{g}(x)=\emptyset$ and $N_{g^{\text{target}}}(x)\neq\emptyset$, $u(g,x)=\frac{1}{2}$; if both are empty set $u(g,x)=1$. The expected payoff is $u(g)=\mathbb{E}_{x\sim\mathcal{D}}\left[u(g,x)\right]=1-\ell^{\textrm{scalar}}(g)$.
@@
Let’s start by focusing on one single point $x$ and its neighborhood. There are $n$ nodes $N_{1}(x)=[\frac{n}{2}]$ and $N_{2}(x)=\{\frac{n}{2}+1,\ldots,n\}$. Consider two graphs—$g_{1}$ with $N_{g_{1}}(x)=N_{1}(x)$ and $g_{2}$ with $N_{g_{2}}(x)=N_{1}(x)\cup N_{2}(x)$. So $N_{g_{1}}(x)$ contains half of the nodes in $N_{g_{2}}(x)$.
In a world characterized by $\beta\in[\frac{1}{8},\frac{2}{3}]$, $N_{g^{\text{target}}}(x)$ is generated in the following random way: Randomly select $\frac{3}{4}\cdot\beta n$ nodes from $N_{1}(x)$ and $\frac{1}{4}\cdot\beta n$ nodes from $N_{2}(x)$. We denote this distribution by $P_{\beta}$. No matter what $\beta$ is, w.p. $\frac{3}{4}$, $v$ is sampled uniformly at random from $N_{1}(x)$ and w.p. $\frac{1}{4}$, $v$ is sampled uniformly at random from $N_{2}(x)$. That is, every node in $N_{1}(x)$ has probability $\frac{3}{2n}$ of being sampled and every node in $N_{2}(x)$ has probability $\frac{1}{2n}$ of being sampled.
@@
$u(g_{1},x)=\frac{|N_{g^{\text{target}}}(x)\cap N_{g_{1}}(x)|}{2|N_{g_{1}}(x)|}+\frac{|N_{g^{\text{target}}}(x)\cap N_{g_{1}}(x)|}{2|N_{g^{\text{target}}}(x)|}=\frac{3/4\cdot\beta n}{n}+\frac{3/4\cdot\beta n}{2\beta n}=\frac{3}{4}\beta+\frac{3}{8}\,,$
@@
$u(g_{2},x)=\frac{|N_{g^{\text{target}}}(x)\cap N_{g_{2}}(x)|}{2|N_{g_{2}}(x)|}+\frac{|N_{g^{\text{target}}}(x)\cap N_{g_{2}}(x)|}{2|N_{g^{\text{target}}}(x)|}=\frac{\beta n}{2n}+\frac{\beta n}{2\beta n}=\frac{1}{2}\beta+\frac{1}{2}\,.$
We make infinite copies of $\{x,N_{1}(x),N_{2}(x)\}$. In each of the copy, $N_{g_{1}}(x)=N_{1}(x)$ and $N_{g_{2}}(x)=N_{1}(x)\cup N_{2}(x)$. For each $x$, we independently sample $N_{g^{\text{target}}}(x)$ from $P_{\beta}$. Let the data distribution over all of such copies of $x$. Then almost surely, there is no repentance in the training data, i.e., there does not exist $i\neq j$ such that $x_{i}=x_{j}$. And for any random sampled test point, w.p. $1$, it has not been sampled in the training set.
Analysis For any unobserved $x\notin\{x_{i}|i\in[m]\}$, let $\alpha_{1}=\frac{|N_{g^{\text{output}}}(x)\cap N_{1}(x)|}{n}$ and $\alpha_{2}=\frac{|N_{g^{\text{output}}}(x)\cap N_{2}(x)|}{n}$. Note that $\alpha_{1},\alpha_{2}$ are in $[0,\frac{1}{2}]$ and are possibly random variables if $\mathcal{A}$ is randomized. Then the expected (over the randomness of $g^{\text{target}}$) payoff of $g^{\text{output}}$ at $x$ is
$\mathbb{E}_{g^{\text{target}}}\left[u(g^{\text{output}},x)\right]=$ $\mathbb{E}_{g^{\text{target}}}\left[\frac{|N_{g^{\text{target}}}(x)\cap N_{g^{\text{output}}}(x)|}{2|N_{g^{\text{output}}}(x)|}+\frac{|N_{g^{\text{target}}}(x)\cap N_{g^{\text{output}}}(x)|}{2|N_{g^{\text{target}}}(x)|}\right]$
@@
Therefore, for any algorithm $\mathcal{A}$, for any $x_{1:m},v_{1:m}$, $\mathbb{E}_{g^{\text{target}}}\left[\ell^{\textrm{scalar}}(g^{\text{output}})\right]$ is worse than $1.05\cdot\min\{\ell^{\textrm{scalar}}(g_{1}),\ell^{\textrm{scalar}}(g_{2})\}$ at either $\beta=\frac{1}{8}$ or $\beta=\frac{2}{3}$. So there exists a target graph such that $\ell^{\textrm{scalar}}(g^{\text{output}})\geq 1.05\cdot\min\{\ell^{\textrm{scalar}}(g_{1}),\ell^{\textrm{scalar}}(g_{2})\}$. ∎
@@ 5.3.2 Proof of Theorem 2
We use the following auxiliary metric between graphs:
@@
For two graphs $g^{\prime},g^{\prime\prime}$ define
@@
For any graph $g$, the scaler loss $\ell^{\textrm{scalar}}(g)=\frac{1}{2}d_{\mathtt{p,r}}(g^{\text{output}},g^{\text{target}})$. In the remainder of this section, we focus on proving Theorem 2. The basic idea is to show that $d_{\mathcal{H}}$ can be used as a surrogate for $d_{\mathtt{p,r}}$. The following lemma plays a crucial role in our proof.
@@
For every pair of graphs $h,k$:
$d_{\mathcal{H}}(h,k)\leq d_{\mathtt{p,r}}(h,k).$
If in addition $h,k\in\mathcal{H}$, we have:
$d_{\mathtt{p,r}}(h,k)\leq 2d_{\mathcal{H}}(h,k).$
@@
$\leq d_{\mathcal{H}}(g,\widehat{g})+\varepsilon/4$ ($g^{\text{output}}\in\arg\min_{g\in\mathcal{H}}d_{\mathcal{H}}(g,\widehat{g})$)
@@
Proof of Lemma 9  For the first inequality, note that both of the distributions $U_{i}^{h}$ and $U_{i}^{k}$ are uniform over their supports and hence $\mathtt{TV}(U_{i}^{h},U_{i}^{k})=\max\{U_{i}^{h}(N_{h}(x_{i})\setminus N_{k}(x_{i})),U_{i}^{k}(N_{k}(x_{i})\setminus N_{h}(x_{i}))\}$. Thus, for every $g^{\prime},g^{\prime\prime}\in\mathcal{H}$:
$\lvert U_{i}^{h}(N_{g^{\prime}}(x_{i})\setminus N_{g^{\prime\prime}}(x_{i}))-U_{i}^{k}(N_{g^{\prime}}(x_{i})\setminus N_{g^{\prime\prime}}(x_{i}))\rvert$
$\leq$ $\mathtt{TV}(U_{i}^{h},U_{i}^{k})$
$\leq$ $U_{i}^{h}(N_{h}(x_{i})\setminus N_{k}(x_{i}))+U_{i}^{k}(N_{k}(x_{i})\setminus N_{h}(x_{i})).$
@@
$d_{\mathcal{H}}(h,k)\leq\frac{1}{m}\sum_{i=1}^{m}\mathtt{TV}(U_{i}^{h},U_{i}^{k})\leq d_{\mathtt{p,r}}(h,k),$
@@
For the second inequality, assume $h,k\in\mathcal{H}$. Thus,
$d_{\mathcal{H}}(h,k)$ $\geq\max\Bigl{\{}\frac{1}{m}\sum_{i=1}^{m}U_{i}^{h}(N_{h}(x_{i})\setminus N_{k}(x_{i})),\frac{1}{m}\sum_{i=1}^{m}U_{i}^{k}(N_{k}(x_{i})\setminus N_{h}(x_{i}))\Bigr{\}}$
$\geq\frac{1}{m}\sum_{i=1}^{m}\frac{U_{i}^{h}(N_{h}(x_{i})\setminus N_{k}(x_{i}))+U_{i}^{k}(N_{k}(x_{i})\setminus N_{h}(x_{i}))}{2}$
$=\frac{1}{2}d_{\mathtt{p,r}}(h,k).$
@@ 5.4 Algorithm and Proofs in the Semi-Realizable Case
$g^{\text{output}}=\operatorname*{arg\,min}_{g\in\mathcal{H}}\sum_{i=1}^{m}\frac{\mathds{1}(v_{i}\in N_{g}(x_{i}))}{n_{g}(x_{i})}\,.$
If there are multiple solutions, we break ties by picking the graph with smallest empirical recall loss.
Proof  For any graph $g$, $\mathds{1}(v_{i}\in N_{g}(x_{i}))$ is an unbiased estimate of the recall $\frac{|N_{g}(x_{i})\cap N_{g^{\text{target}}}(x_{i})|}{n_{g^{\text{target}}}(x_{i})}$. Thus, $\frac{\mathds{1}(v_{i}\in N_{g}(x_{i}))}{n_{g}(x_{i})}$ is an unbiased estimate of $\frac{|N_{g}(x_{i})\cap N_{g^{\text{target}}}(x_{i})|}{n_{g}(x_{i})\cdot n_{g^{\text{target}}}(x_{i})}$. Since $g^{\prime}$ has zero precision loss, $\frac{|N_{g}(x_{i})\cap N_{g^{\text{target}}}(x_{i})|}{n_{g}(x_{i})}=1$ almost everywhere. Thus, we have
$\left|\frac{1}{m}\sum_{i=1}^{m}\frac{\mathds{1}(v_{i}\in N_{g}(x_{i}))}{n_{g}(x_{i})}-\mathbb{E}\left[\frac{1}{n_{g^{\text{target}}}(x)}\right]\right|\leq\sqrt{\frac{\log(|\mathcal{H}|)+\log(1/\delta)}{m}}\,,$
for all $g\in\mathcal{H}$. Then if $\Delta_{\mathcal{D}}>0$, we need $\frac{1}{\Delta_{\mathcal{D}}^{2}}$ samples to separate $g^{\prime}$ from other graphs in the hypothesis class. ∎
@@
Proof  Let’s start by focusing on one single point $x$. Let $N=\{v_{1},\ldots,v_{n}\}$ for some $n\gg m$. Let $N_{g_{1}}(x)=\{v_{1}\}$ and $N_{g_{2}}(x)=\{v_{2}\}$. In world I, $N_{g^{\text{target}}}(x)$ is generated in the following way.
  • w.p. $\frac{1}{2}$, $N_{g^{\text{target}}}(x)=N\setminus\{v_{2}\}$.
  • w.p. $\frac{1}{2}$, $N_{g^{\text{target}}}(x)=\{v_{1},v_{2}\}$.
@@
  • w.p. $\frac{1}{2}$, $N_{g^{\text{target}}}(x)=N\setminus\{v_{1}\}$.
  • w.p. $\frac{1}{2}$, $N_{g^{\text{target}}}(x)=\{v_{1},v_{2}\}$.
We make infinite independent copies of $(x,N)$ and let $\mathcal{D}$ to be the uniform distribution over such $x$’s. Hence, in world I, $\ell^{\textrm{precision}}(g_{1})=0$ and $\ell^{\textrm{recall}}(g_{1})=\frac{3}{4}-\frac{1}{2n}$; $\ell^{\textrm{precision}}(g_{2})=\frac{1}{2}$ and $\ell^{\textrm{recall}}(g_{2})=\frac{3}{4}$. When $n\rightarrow\infty$, we can’t distinguish between two worlds. In order to achieve $\ell^{\textrm{precision}}(g^{\text{output}})=0$ in both worlds, we need to make $N_{g^{\text{output}}}(x)=\emptyset$ for almost every $x$. Then the recall loss would be $1$. ∎

### v2

@@ 5.3 Surrogate Loss Method in Both Realizable and Agnostic Cases
Again we focus empirical precision and recall losses minimization. Let $x_{1},\ldots,x_{m}\in\mathcal{X}$ denote a sequence of inputs. For each hypothesis $g$, let $U_{i}^{g}$ denote the uniform distribution over the output $g(x_{i})$. For any pair of hypotheses $g^{\prime},g^{\prime\prime}$, define the following:
$\mathtt{precision.loss}(g^{\prime}\mid g^{\prime\prime})$ $=\mathtt{recall.loss}(g^{\prime\prime}\mid g^{\prime}):=\frac{1}{m}\sum_{i=1}^{m}U_{i}^{g^{\prime}}(g^{\prime}(x_{i})\setminus g^{\prime\prime}(x_{i}))\,.$
Here $\mathtt{precision.loss}(g^{\prime}\mid g^{\prime\prime})$ is the precision loss of hypothesis $g^{\prime}$ when the target hypothesis is $g^{\prime\prime}$ and $\mathtt{recall.loss}(g^{\prime\prime}\mid g^{\prime})$ is the recall loss of hypothesis $g^{\prime\prime}$ when the target hypothesis is $g^{\prime}$. Thus, the goal is to output a hypothesis $g$ with small $\mathtt{precision.loss}(g\mid g^{\text{target}})$ and $\mathtt{recall.loss}(g\mid g^{\text{target}})$.
@@
For a hypothesis $g$ define a vector $v_{g}:\mathcal{H}\times\mathcal{H}\to[0,1]$ by
$v_{g}(g^{\prime},g^{\prime\prime})=\frac{1}{m}\sum_{i=1}^{m}U_{i}^{g}(g^{\prime}(x_{i})\setminus g^{\prime\prime}(x_{i})).$
Define a metric $d_{\mathcal{H}}$ between hypotheses by $d_{\mathcal{H}}(g_{1},g_{2})=\|v_{g_{1}}-v_{g_{2}}\|_{\infty}$.
Let $\widehat{g}$ be the observed (empirical) hypothesis; i.e. the hypothesis which outputs $\{v_{i}\}$ at $x_{i}$ for all $(x_{i},v_{i})$ in the training set and outputs the empty set for all unobserved input. A standard union bound argument yields:
@@
Let $g^{\text{target}}$ denote the true hypothesis (i.e. the data is generated from $g^{\text{target}}$). Then, with probability at least $1-\delta$:
@@
Algorithm (realizable case): Let $\varepsilon$ denote the desired error. Output a hypothesis $g^{\text{output}}\in\mathcal{H}$ such that
@@
Notice that Item 1 corresponds to the first principle for discarding suboptimal hypotheses described earlier in this section, while Item 2 corresponds to the second principle.
Algorithm (agnostic case): output a hypothesis $g^{\text{output}}\in\mathcal{H}$ such that
@@
Let $g^{\text{target}}$ denote the target hypothesis. Then, for
@@
the agnostic-case algorithm outputs a hypothesis $g^{\text{output}}$ such that with probability at least $1-\delta$,
@@ 5.3.1 Proof of Theorem 3
Proof of Theorem 3 For simplicity, we adopt payoffs instead of losses here. The payoff of hypothesis $g$ at $x$ is
$u(g,x)=\frac{|g^{\text{target}}(x)\cap g(x)|}{2|g(x)|}+\frac{|g^{\text{target}}(x)\cap g(x)|}{2|g^{\text{target}}(x)|}\,.$
If $g(x)=\emptyset$ and $g^{\text{target}}(x)\neq\emptyset$, $u(g,x)=\frac{1}{2}$; if both are empty set $u(g,x)=1$. The expected payoff is $u(g)=\mathbb{E}_{x\sim\mathcal{D}}\left[u(g,x)\right]=1-\ell^{\textrm{scalar}}(g)$.
@@
[…]
In a world characterized by $\beta\in[\frac{1}{8},\frac{2}{3}]$, $g^{\text{target}}(x)$ is generated in the following random way: Randomly select $\frac{3}{4}\cdot\beta n$ items from $N_{1}(x)$ and $\frac{1}{4}\cdot\beta n$ items from $N_{2}(x)$. We denote this distribution by $P_{\beta}$. No matter what $\beta$ is, w.p. $\frac{3}{4}$, $v$ is sampled uniformly at random from $N_{1}(x)$ and w.p. $\frac{1}{4}$, $v$ is sampled uniformly at random from $N_{2}(x)$. That is, every item in $N_{1}(x)$ has probability $\frac{3}{2n}$ of being sampled and every item in $N_{2}(x)$ has probability $\frac{1}{2n}$ of being sampled.
@@
$u(g_{1},x)=\frac{|g^{\text{target}}(x)\cap g_{1}(x)|}{2|g_{1}(x)|}+\frac{|g^{\text{target}}(x)\cap g_{1}(x)|}{2|g^{\text{target}}(x)|}=\frac{3/4\cdot\beta n}{n}+\frac{3/4\cdot\beta n}{2\beta n}=\frac{3}{4}\beta+\frac{3}{8}\,$,
@@
$u(g_{2},x)=\frac{|g^{\text{target}}(x)\cap g_{2}(x)|}{2|g_{2}(x)|}+\frac{|g^{\text{target}}(x)\cap g_{2}(x)|}{2|g^{\text{target}}(x)|}=\frac{\beta n}{2n}+\frac{\beta n}{2\beta n}=\frac{1}{2}\beta+\frac{1}{2}\,.$
We make infinite copies of $\{x,N_{1}(x),N_{2}(x)\}$. In each of the copy, $g_{1}(x)=N_{1}(x)$ and $g_{2}(x)=N_{1}(x)\cup N_{2}(x)$. For each $x$, we independently sample $g^{\text{target}}(x)$ from $P_{\beta}$. Let the data distribution over all of such copies of $x$. Then almost surely, there is no repentance in the training data, i.e., there does not exist $i\neq j$ such that $x_{i}=x_{j}$. And for any random sampled test point, w.p. $1$, it has not been sampled in the training set.
Analysis For any unobserved $x\notin\{x_{i}|i\in[m]\}$, let $\alpha_{1}=\frac{|g^{\text{output}}(x)\cap N_{1}(x)|}{n}$ and $\alpha_{2}=\frac{|g^{\text{output}}(x)\cap N_{2}(x)|}{n}$. Note that $\alpha_{1},\alpha_{2}$ are in $0,\frac{1}{2}]$ and are possibly random variables if $\mathcal{A}$ is randomized. Then the expected (over the randomness of $g^{\text{target}}$) payoff of $g^{\text{output}}$ at $x$ is
$\mathbb{E}_{g^{\text{target}}}\left[u(g^{\text{output}},x)\right]=$ $\mathbb{E}_{g^{\text{target}}}\left[\frac{|g^{\text{target}}(x)\cap g^{\text{output}}(x)|}{2|g^{\text{output}}(x)|}+\frac{|g^{\text{target}}(x)\cap g^{\text{output}}(x)|}{2|g^{\text{target}}(x)|}\right]$
@@
Therefore, for any algorithm $\mathcal{A}$, for any $x_{1:m},v_{1:m}$, $\mathbb{E}_{g^{\text{target}}}\left[\ell^{\textrm{scalar}}(g^{\text{output}})\right]$ is worse than $1.05\cdot\min\{\ell^{\textrm{scalar}}(g_{1}),\ell^{\textrm{scalar}}(g_{2})\}$ at either $\beta=\frac{1}{8}$ or $\beta=\frac{2}{3}$. So there exists a target hypothesis such that $\ell^{\textrm{scalar}}(g^{\text{output}})\geq 1.05\cdot\min\{\ell^{\textrm{scalar}}(g_{1}),\ell^{\textrm{scalar}}(g_{2})\}$. ∎
@@ 5.3.2 Proof of Theorem 2
We use the following auxiliary metric between hypotheses:
@@
For two hypotheses $g^{\prime},g^{\prime\prime}$ define
@@
For any hypothesis $g$, the scaler loss $\ell^{\textrm{scalar}}(g)=\frac{1}{2}d_{\mathtt{p,r}}(g^{\text{output}},g^{\text{target}})$. In the remainder of this section, we focus on proving Theorem 2. The basic idea is to show that $d_{\mathcal{H}}$ can be used as a surrogate for $d_{\mathtt{p,r}}$. The following lemma plays a crucial role in our proof.
@@
For every pair of hypotheses $g_{1},g_{2}$:
$d_{\mathcal{H}}(g_{1},g_{2})\leq d_{\mathtt{p,r}}(g_{1},g_{2}).$
If in addition $g_{1},g_{2}\in\mathcal{H}$, we have:
$d_{\mathtt{p,r}}(g_{1},g_{2})\leq 2d_{\mathcal{H}}(g_{1},g_{2}).$
@@
[…]
@@
Proof of Lemma 9 For the first inequality, note that both of the distributions $U_{i}^{g_{1}}$ and $U_{i}^{g_{2}}$ are uniform over their supports and hence $\mathtt{TV}(U_{i}^{g_{1}},U_{i}^{g_{2}})=\max\{U_{i}^{g_{1}}(g_{1}(x_{i})\setminus g_{2}(x_{i})),U_{i}^{g_{2}}(g_{2}(x_{i})\setminus g_{1}(x_{i}))\}$. Thus, for every $g^{\prime},g^{\prime\prime}\in\mathcal{H}$:
$\lvert U_{i}^{g_{1}}(g^{\prime}(x_{i})\setminus g^{\prime\prime}(x_{i}))-U_{i}^{g_{2}}(g^{\prime}(x_{i})\setminus g^{\prime\prime}(x_{i}))\rvert$
$\leq$ $\mathtt{TV}(U_{i}^{g_{1}},U_{i}^{g_{2}})$
$\leq$ $U_{i}^{g_{1}}(g_{1}(x_{i})\setminus g_{2}(x_{i}))+U_{i}^{g_{2}}(g_{2}(x_{i})\setminus g_{1}(x_{i})).$
@@
$d_{\mathcal{H}}(g_{1},g_{2})\leq\frac{1}{m}\sum_{i=1}^{m}\mathtt{TV}(U_{i}^{g_{1}},U_{i}^{g_{2}})\leq d_{\mathtt{p,r}}(g_{1},g_{2}),$
@@
For the second inequality, assume $g_{1},g_{2}\in\mathcal{H}$. Thus,
$d_{\mathcal{H}}(g_{1},g_{2})$ $\geq\max\Bigl\{\frac{1}{m}\sum_{i=1}^{m}U_{i}^{g_{1}}(g_{1}(x_{i})\setminus g_{2}(x_{i})),\frac{1}{m}\sum_{i=1}^{m}U_{i}^{g_{2}}(g_{2}(x_{i})\setminus g_{1}(x_{i}))\Bigr\}$
$\geq\frac{1}{m}\sum_{i=1}^{m}\frac{U_{i}^{g_{1}}(g_{1}(x_{i})\setminus g_{2}(x_{i}))+U_{i}^{g_{2}}(g_{2}(x_{i})\setminus g_{1}(x_{i}))}{2}$
$=\frac{1}{2}d_{\mathtt{p,r}}(g_{1},g_{2}).$
@@ 5.4 Algorithm and Proofs in the Semi-Realizable Case
$g^{\text{output}}=\operatorname*{arg\,min}_{g\in\mathcal{H}}\sum_{i=1}^{m}\frac{\mathds{1}(v_{i}\in g(x_{i}))}{n_{g}(x_{i})}\,.$
If there are multiple solutions, we break ties by picking the hypothesis with smallest empirical recall loss.
Proof of Theorem 3 For any hypothesis $g$, $\mathds{1}(v_{i}\in g(x_{i}))$ is an unbiased estimate of the recall $\frac{|g(x_{i})\cap g^{\text{target}}(x_{i})|}{n_{g^{\text{target}}}(x_{i})}$. Thus, $\frac{\mathds{1}(v_{i}\in g(x_{i}))}{n_{g}(x_{i})}$ is an unbiased estimate of $\frac{|g(x_{i})\cap g^{\text{target}}(x_{i})|}{n_{g}(x_{i})\cdot n_{g^{\text{target}}}(x_{i})}$.
Since $g^{\prime}$ has zero precision loss, $\frac{|g(x_{i})\cap g^{\text{target}}(x_{i})|}{n_{g}(x_{i})}=1$ almost everywhere. Thus, we have
$\left|\frac{1}{m}\sum_{i=1}^{m}\frac{\mathds{1}(v_{i}\in g(x_{i}))}{n_{g}(x_{i})}-\mathbb{E}\left[\frac{1}{n_{g^{\text{target}}}(x)}\right]\right|\leq\sqrt{\frac{\log(|\mathcal{H}|)+\log(1/\delta)}{m}}\,$,
for all $g\in\mathcal{H}$. Then if $\Delta_{\mathcal{D}}>0$, we need $\frac{1}{\Delta_{\mathcal{D}}^{2}}$ samples to separate $g^{\prime}$ from other hypotheses in the hypothesis class. ∎
@@
Proof of Theorem 3 Let’s start by focusing on one single input $x$. Let $N=\{v_{1},\ldots,v_{n}\}$ for some $n\gg m$. Let $g_{1}(x)=\{v_{1}\}$ and $g_{2}(x)=\{v_{2}\}$. In world I, $g^{\text{target}}(x)$ is generated in the following way.
  • w.p. $\frac{1}{2}$, $g^{\text{target}}(x)=N\setminus\{v_{2}\}$.
  • w.p. $\frac{1}{2}$, $g^{\text{target}}(x)=\{v_{1},v_{2}\}$.
@@
  • w.p. $\frac{1}{2}$, $g^{\text{target}}(x)=N\setminus\{v_{1}\}$.
  • w.p. $\frac{1}{2}$, $g^{\text{target}}(x)=\{v_{1},v_{2}\}$.
We make infinite independent copies of $(x,N)$ and let $\mathcal{D}$ to be the uniform distribution over such $x$’s. Hence, in world I, $\ell^{\textrm{precision}}(g_{1})=0$ and $\ell^{\textrm{recall}}(g_{1})=\frac{3}{4}-\frac{1}{2n}$; $\ell^{\textrm{precision}}(g_{2})=\frac{1}{2}$ and $\ell^{\textrm{recall}}(g_{2})=\frac{3}{4}$. When $n\rightarrow\infty$, we can’t distinguish between two worlds. In order to achieve $\ell^{\textrm{precision}}(g^{\text{output}})=0$ in both worlds, we need to make $g^{\text{output}}(x)=\emptyset$ for almost every $x$. Then the recall loss would be $1$. ∎